OpenStudy (anonymous):

which one of the following numbers will appear farthest to the right on the number line? A.8/3 B pi 3.14 C.square root of 5 D.2.32

OpenStudy (anonymous):

\[\frac{8}{3} = 2.66666...\]\[\pi=3.14\]\[\sqrt{5} = 2.236\]\[2.32\] The number that appears farthest to the right will be the number of greatest value.

OpenStudy (anonymous):

so B

OpenStudy (anonymous):

Yes, sir.

OpenStudy (anonymous):

thanks